A sample of cupcakes contains 124 pink ones with 3 white ones. How many white ones can you expect in a sample of 1270 total cupcakes?
Please show all steps
I get a decimal and believe it is incorrect, thank you for the assistance

Answer by stanbon(75887)

A sample of cupcakes contains 124 pink ones with 3 white ones. How many white ones can you expect in a sample of 1270 total cupcakes?
----
Solve the proportion:
x/1270 = 3/(3+124)
----
x = 1270[3/(127)
---
x = 3810/127
x = 30